Parking Meter Theft Listed below are amounts (in millions of dollars) collected from parking meters by Brinks and others in New York City during similar time periods. A larger data set was used to convict five Brinks employees of grand larceny. The data were provided by the attorney for New York City, and they are listed on the DASL Web site.

Use s=0.178 for Brinks employees            

Find the coefficient of variation for the samples.

Collection contractor was Brinks:       1.3 1.5 1.3 1.5 1.4 1.7 1.8 1.7 1.7 1.6

Group of answer choices

a) C.V. = -11.5%

b) C.V. = 10.5%

c) C.V. = 21.5%

d) C.V. = 11.5%
Answer #1

Solution:

Option (d) is correct.

Explanation:

Mean = 1.55

Standard deviation = 0.1780

C.V. = \frac{\sigma}{\bar{X}}\times 100

C.V. =\frac{0.1780}{1.55}\times 100

C.V. = 11.48% = 11.5%

So, Option (d) is correct.

R PROGRAM:

x=c( 1.3 ,1.5 ,1.3 ,1.5, 1.4 ,1.7 ,1.8 ,1.7 ,1.7 ,1.6)
x
sd(x)
mean(x)

PLEASE RATING THE ANSWER. THANK YOU....!!